Felipe is a teenager who is 10 years older than his sister Irene. In 6 years, Felipe will be twice as old as his sister. How old is Felipe now?
Can you please show us how to set up this equation, and the process to solve it?

Answer by askmemath(368)

Let Irene's age = X
Since Felipe is older by 10 years , his age = X+10
In 6 years
Irene's age = X+6
Felipe's age = X+10 +6 = X+16
Felipe will be twice as old in 6 yrs
X+16 = 2(X+6)
X+16 = 2X+12
Subtracting X+12 on both sides
X = 4 (irene's age)
Felipe's age = 4+10 = 14